Between two A and B there are 12 intermediate station how many ways a train can stop at three 
Intermediate station such that no two stopea are 
Conjugative 

Grade:12th pass

1 Answers

Rajat
4 years ago
Total number of ways of choosing three stops among 12 intermediate stations= 12C3 = 220.
Number of ways in which 2 stops are consecutive= 11
and number ways of choosing the third stop = 10 
Thus number of ways two or more stops are consecutive = 11×10=110
Thus required number of ways = 220-110= 110